# A Survey on Human-AI Teaming with Large Pre-Trained Models

### v1

@@ 3.3.3 Evaluation the Effectiveness of Joint Human-AI Systems
Evaluating the efficacy of human-AI systems extends beyond technical performance, encompassing a holistic approach that includes technical, social, and psychological dimensions of human-AI collaboration. The study by Henry et al. [63] emphasizes the importance of user trust and system usability, suggesting that evaluation frameworks should assess how systems support user autonomy and are endorsed by domain experts to facilitate adoption. Similarly, Liu et al. [94] highlight the need to understand perceptual differences between humans and AI, advocating for evaluations that measure how these systems leverage the complementary strengths of human and machine intelligence to improve joint decision-making processes. Munyaka et al. [112] delve into the impact of decision-making styles and AI identity disclosure on team dynamics and efficacy, underlining the significance of evaluating social dynamics and collaboration patterns within human-AI teams. Furthermore, Memmert et al. [105] explore the integration of AI in creative processes, pointing to the necessity of assessing both the cognitive contributions of AI and its effects on group dynamics, such as the potential for free-riding.
@@
In terms of large models, Chang et al. [22] present a comprehensive theoretical framework for evaluating LLMs. This framework spans a wide array of metrics and scenarios, such as natural language processing, reasoning capabilities, and robustness. This study underscores the necessity of multifaceted evaluation frameworks that extend beyond task performance to include considerations like ethical implications and bias, which are crucial in human-AI systems. On the other hand, Veselovsky et al.[157] adopt a more empirical approach by examining the practical use of LLMs in crowd work. Their evaluation method involves observing and adapting the behavior of crowd workers as they interact with LLMs, offering a practical perspective on Human-AI interaction. These studies collectively suggest that a comprehensive evaluation of human-AI systems should not only focus on accuracy or efficiency but also consider factors like trust, usability, collaboration dynamics, and the psychological impact on users, to ensure the development of systems that are effectively integrated into human workflows and decision-making processes.
4 Safe, Secure and Trustworthy AI
@@
4.1 Algorithmic Bias and Fairness
@@
Algorithmic fairness is a cornerstone of contemporary AI, demanding the development and deployment of unbiased and non-discriminatory algorithms. This pursuit, though challenging due to the complexities in defining fairness, mitigating data biases, and ensuring algorithm transparency, has become a focal point for researchers and organizations [54]. Fairness in algorithms is inherently complex. For instance, biases in the training data can lead to skewed outputs, and inherent biases in algorithmic design can persist even with unbiased data. These issues are further complicated by the often opaque nature of algorithms and the lack of established accountability measures [143]. Several strategies have been proposed to address these challenges. Key among them are ensuring data representativeness and employing algorithms known for their fairness. Pre-deployment testing for biases, enhancing transparency, and establishing robust accountability mechanisms are equally crucial [33]. Chhibber et al’s [25] research sheds light on the role of user interaction in shaping algorithmic outcomes. It demonstrates how data quality, user engagement, and system localization impact fairness in machine learning systems. Additionally, Morrison et al.’s [111] study underscores the value of causal explanations in AI, arguing that they foster trust, facilitate error correction, and support informed decision-making, thereby contributing to the enhancement of algorithmic fairness.
@@
The incorporation of Large Language Models (LLMs) into various applications has intensified the need to address algorithmic bias and fairness. Kirk et al. [81] have shown that biases in large language models, particularly those available ’out-of-the-box’ like GPT-2, can lead to stereotypical and less diverse occupational associations for women, especially when intersecting with other categories like religion or ethnicity. Meanwhile, Huang et al. [66] reveal that a significant portion of code functions generated by LLMs can contain biases related to sensitive attributes such as age and gender, raising concerns about the fairness and integrity of software applications reliant on these models. Meyer et al. [106] explore the paradigm shift brought by conversational LLMs, discussing their potential to enhance academic work while cautioning against their inherent biases and accuracy issues. Kotek et al. [83] further investigate gender stereotypes within LLMs, demonstrating that these models are significantly more likely to align occupations with stereotypical gender roles, sometimes amplifying biases beyond societal norms or official statistics.
@@
Recent studies are also working on diverse strategies for mitigating bias in LLMs, emphasizing the importance of fairness across their applications. Li et al. [92] review fairness strategies for both medium and large-sized LLMs, highlighting the need for tailored debiasing methods based on model size and training paradigms. Gallegos et al. [50] offer a detailed survey on bias evaluation and mitigation, categorizing techniques across model development stages and introducing taxonomies for bias metrics and mitigation methods. Ohi et al. [119] propose a novel approach to counteract likelihood bias in LLM evaluations using few-shot examples for in-context learning, showing marked improvements in fairness. Bi et al. [10] focus on group fairness in social media, presenting a hierarchical schema for bias evaluation and a chain-of-thought mitigation method, GF-Think, to reduce bias. These contributions underline possible approaches to bias mitigation, aiming for ethical and equitable use of LLMs.
4.2 Worker Autonomy and Well Being
@@
Chhibber et al. [25] demonstrate how teachable conversational agents in crowd-based applications can empower workers, enhancing control, personalized learning, and ownership, thereby boosting job satisfaction. Konstantis et al. [82] emphasize transparency and fairness in crowdsourcing platforms, advocating for clear decision-making and respectful treatment to improve worker experience. Furthermore, Pal et al. [123] highlight the design of LLMs focused on worker well-being and autonomy, avoiding stress-inducing features and promoting engagement and learning.
Despotovic et al. [37] highlight a trend in job seekers preferring roles that incorporate advanced AI, like ChatGPT, aligning with their identities and technological interests. This shift towards AI-interactive jobs indicates a broader preference for technologically engaging roles, impacting job satisfaction and autonomy. Complementarily, Wang et al. [162] explore the paradox of LLMs in the job market, noting their role in both creating new opportunities and obsoleting certain jobs. This dual effect is key in assessing worker autonomy and well-being, as LLM-driven automation fosters more creative roles and autonomy but also raises concerns about job security and satisfaction. Integrating insights from recent research, it’s evident that while generative AI has the potential to automate menial tasks and foster more creative roles, it also poses challenges to worker autonomy, necessitating a balanced approach to AI deployment in the workplace [170].
4.3 Effect on Wages and Jobs
@@
Hemmer et al. [62] suggest AI could enhance productivity by up to 40%, potentially leading to higher wages. In contrast, Chowdhury et al. [28] point to AI’s role in both creating and displacing jobs, with an estimated 1.7 million new jobs and significant displacements in the U.S. by 2029. This highlights the necessity for proactive support for those affected by AI-driven changes. Chhibber et al. [25] discuss AI’s impact on the gig economy, noting potential job losses in areas like customer service, balanced by new opportunities in AI-related fields. Pal et al. [123] foresee a growing LLM industry, suggesting increased job creation in AI R&D. Conversely, Wang et al. [162] anticipate the displacement of up to 800 million jobs globally by 2030, advocating for protective legislation and equitable AI benefit distribution. Ashktorab et al. [3] examine Human-AI collaboration in the gaming industry, where AI’s role in automating tasks like game testing could lead to job enrichment in creative roles.
@@
The integration of Large Language Models has diverse implications for the job market. Despotovic et al.’s [37] research reveals that job seekers show a preference for roles that resonate with their personal identity and involve frequent interaction with technologies like ChatGPT. This underscores a shift in job market dynamics and the necessity for businesses to adapt their recruitment strategies. Wang et al. [162] discuss the paradoxical effect of LLMs, highlighting both job creation and displacement across various sectors. This includes an examination of LLMs’ influence on routine tasks, specialized professions, and broader economic and societal aspects. Walkowiak et al. [158] focus on the Australian workforce, quantifying the exposure to GenAI risks and potential job transformations, addressing privacy, cybersecurity, and ethical issues.
4.4 Data Privacy and Security
@@
In the realm of human-AI collaboration, trust, particularly in data privacy and security, is foundational. Ezer et al. [43] emphasize the critical role of trust engineering, advocating for algorithmic transparency and dynamic management to establish mutual trust in complex, information-imperfect environments. Yin et al. [176] further this by proposing a privacy-focused, human-centric approach, leveraging logistic regression and differential privacy to strike a balance between privacy and utility, achieving notable classification accuracy. In medical imaging, Kaissis et al. [74] highlight the imperative for secure, privacy-preserving AI through federated learning, addressing the privacy-utilization dichotomy to advance patient care. The synergy of AI and cybersecurity, as explored by Admin et al. [136], showcases AI’s potential to fortify cyber defenses, underscoring the role of machine and deep learning in countering complex cyber threats and safeguarding data integrity. Lepri et al. [89] advocate for a human-centric AI ethos, stressing the importance of privacy, accountability, and ethical decision-making in building secure, trustworthy AI systems.
@@
Regulatory and privacy challenges for Large Generative AI Models (LGAIMs), including ChatGPT and GPT-4, demand targeted legal and technical solutions for safe use. Hacker et al. [57] recommend a regulatory framework for LGAIMs focusing on transparency, risk management, and equitable treatment, particularly for high-risk uses. Ullah et al. [153] propose PrivChatGPT, a model incorporating differential privacy for LLMs, highlighting the balance between privacy and utility. Gupta et al. [56] discuss cybersecurity risks with GenAI, emphasizing the need for ethical practices and secure defenses against potential misuse. De Angelis et al. [36] identify LLMs as vectors for misinformation, advocating for policies to combat the "AI-driven infodemic" in public health. Thapa and Adhikari [149] balance LLMs’ benefits against misinformation risks in biomedical research, calling for strict validation. Sebastian [137] focuses on privacy in AI chatbots, suggesting federated learning and data minimization to protect user information, reflecting increasing privacy concerns among LLM users.
4.5 Trustworthy AI and Accountability
@@
Initially, conventional methods in trustworthy AI emphasized the importance of ethical principles, human oversight, and transparency. Hou et al. [64] discuss these principles, stressing the need for secure systems that identify risks while remaining under human control. Pflanzer et al. [126] introduce the Agent-Deed-Consequence model as a traditional ethical algorithm framework, emphasizing human oversight in AI ethics. Caldwell et al. [18] highlight factors influencing trust in AI in cooperative environments, such as multiplayer gaming, underscoring the importance of environmental factors and peer influence in conventional trust-building.
@@
The advent of LLMs introduces new dimensions to these conventional methods. In healthcare, a case study using LLMs [162] showcases the balance between patient engagement and AI over-reliance risks, underscoring the advancements in data sensitivity and patient communication. Kim et al. [79] demonstrate how LLMs contribute to managing uncertainty in human-AI collaboration, proposing a trial-to-trial approach to effectively communicate and estimate uncertainty. Furthermore, LLMs influence the formation of mental models, as discussed by Bansal et al. [6]. They highlight LLMs’ role in enhancing predictability and understanding of AI’s error boundaries, crucial for developing accurate mental models for better human-AI collaboration.
4.6 Law and Public Policy
@@
Traditionally, AI legal and policy frameworks have focused on ensuring transparency, accountability, fairness, and bias mitigation. Pioneering work by Jobin et al. [70], which analyzed AI guidelines globally, underscores the universal emphasis on ethical standards, transparency, and accountability. Similarly, research by Cath et al. [19] highlights the critical need for transparent and accountable AI, especially in decision-making processes. Stahl et al.[143] propose six guidelines for AI regulation: Transparency, Accountability, Bias Mitigation, Human Oversight, Data Protection, and Public Education.
@@
The integration of Large Language Models (LLMs) into the AI landscape underscores the need for updated legal frameworks to address new challenges in data use and AI-generated content regulation. Bender et al. [12] examine the ethical implications of LLMs, focusing on data usage and biases. An accredited law partner [159] specializing in technology and intellectual property discusses U.S. court cases like GitHub Copilot’s use of open-source code, highlighting the legal intricacies of data attribution and compliance. This situation prompts a reevaluation of the datasets used for AI training and the adaptation of laws, particularly concerning copyrights and patents. Ekenobi et al. [148] explore the legal impacts of deploying LLMs, with a focus on data protection and privacy. They commend ChatGPT’s privacy measures as a step towards mitigating potential risks. Additionally, Marcos and Pullin’s article [101] talks about the EU’s GDPR challenges, emphasizing the balance between technological growth and privacy rights. This necessitates robust regulatory frameworks that ensure LLM compliance with strict data protection standards, highlighting the importance of new laws in safeguarding privacy while fostering innovation.
5 Applications

### v2

@@ 3.3.3. Evaluating the Effectiveness of Joint Human-AI Systems
Evaluating the efficacy of human-AI systems extends beyond technical performance, encompassing a holistic approach that includes technical, social, and psychological dimensions of human-AI collaboration. The study by Henry et al. (Henry et al., 2022) emphasizes the importance of user trust and system usability, suggesting that evaluation frameworks should assess how systems support user autonomy and are endorsed by domain experts to facilitate adoption. Similarly, Liu et al. (Liu et al., 2023b) highlight the need to understand perceptual differences between humans and AI, advocating for evaluations that measure how these systems leverage the complementary strengths of human and machine intelligence to improve joint decision-making processes. Munyaka et al. (Munyaka et al., 2023) delve into the impact of decision-making styles and AI identity disclosure on team dynamics and efficacy, underlining the significance of evaluating social dynamics and collaboration patterns within human-AI teams. Furthermore, Memmert et al. (Memmert and Tavanapour, 2023) explore the integration of AI in creative processes, pointing to the necessity of assessing both the cognitive contributions of AI and its effects on group dynamics, such as the potential for free-riding.
@@
In terms of large models, Chang et al. (Chang et al., 2023) present a comprehensive theoretical framework for evaluating LLMs. This framework spans a wide array of metrics and scenarios, such as natural language processing, reasoning capabilities, and robustness. This study highlights the necessity of multifaceted evaluation frameworks that extend beyond task performance to include considerations like ethical implications and bias, which are crucial in human-AI systems. On the other hand, Veselovsky et al.(Veselovsky et al., 2023) adopt a more empirical approach by examining the practical use of LLMs in crowd work. Their evaluation method involves observing and adapting the behavior of crowd workers as they interact with LLMs, offering a practical perspective on Human-AI interaction. These studies collectively suggest that a comprehensive evaluation of human-AI systems should not only focus on accuracy or efficiency but also consider factors like trust, usability, collaboration dynamics, and the psychological impact on users, to ensure the development of systems that are effectively integrated into human workflows and decision-making processes.
4. Safe, Secure and Trustworthy AI
@@
Figure 6. Large Language Models often exhibit gender bias due to primarily being trained on large datasets of human language that reflect existing societal biases. Kotek et al. (Kotek et al., 2023) test this by probing LLMs to answer permutations of complex questions, as shown above. The answers and their explanation help the authors to quantify the bias.
4.1. Algorithmic Bias and Fairness
@@
Algorithmic fairness is a cornerstone of contemporary AI, demanding the development and deployment of unbiased and non-discriminatory algorithms. This pursuit, though challenging due to the complexities in defining fairness, mitigating data biases, and ensuring algorithm transparency, has become a focal point for researchers and organizations (Green and Chen, 2019). Fairness in algorithms is inherently complex. For instance, biases in the training data can lead to skewed outputs, and inherent biases in algorithmic design can persist even with unbiased data. These issues are further complicated by the often opaque nature of algorithms and the lack of established accountability measures (Stahl et al., 2021). Several strategies have been proposed to address these challenges. Key among them are ensuring data representativeness and employing algorithms known for their fairness. Pre-deployment testing for biases, enhancing transparency, and establishing robust accountability mechanisms are equally crucial (Conlon et al., 2023). Chhibber et al’s (Chhibber et al., 2022) research sheds light on the role of user interaction in shaping algorithmic outcomes. It demonstrates how data quality, user engagement, and system localization impact fairness in machine learning systems. Additionally, Morrison et al.’s (Morrison et al., 2023) study discusses the value of causal explanations in AI, arguing that they foster trust, facilitate error correction, and support informed decision-making, thereby contributing to the enhancement of algorithmic fairness.
@@
The incorporation of Large Language Models (LLMs) into various applications has intensified the need to address algorithmic bias and fairness. Kirk et al. (Kirk et al., 2021) have shown that biases in large language models, particularly those available ‘out-of-the-box’ like GPT-2, can lead to stereotypical and less diverse occupational associations for women, especially when intersecting with other categories like religion or ethnicity. Meanwhile, Huang et al. (Huang et al., 2023) reveal that a significant portion of code functions generated by LLMs can contain biases related to sensitive attributes such as age and gender, raising concerns about the fairness and integrity of software applications reliant on these models. Kotek et al. (Kotek et al., 2023) further investigate gender stereotypes within LLMs (Fig. 6), demonstrating that these models are significantly more likely to align occupations with stereotypical gender roles, sometimes amplifying biases beyond societal norms or official statistics. They find out that, on average, these LLMs are 6.8 times more likely to choose a stereotypical female answer when a female pronoun is present and 3.4 times for male answers and pronouns. Further, Meyer et al. (Meyer et al., 2023) explore the paradigm shift brought by conversational LLMs, discussing their potential to enhance academic work while cautioning against their inherent biases and accuracy issues.
@@
Recent studies are also working on diverse strategies for mitigating bias in LLMs, emphasizing the importance of fairness across their applications. Li et al. (Li et al., 2023) review fairness strategies for both medium and large-sized LLMs, highlighting the need for tailored debiasing methods based on model size and training paradigms. Gallegos et al. (Gallegos et al., 2023) offer a detailed survey on bias evaluation and mitigation, categorizing techniques across model development stages and introducing taxonomies for bias metrics and mitigation methods. Ohi et al. (Ohi et al., 2024) propose a novel approach to counteract likelihood bias in LLM evaluations using few-shot examples for in-context learning, showing marked improvements in fairness. Bi et al. (Bi et al., 2023) focus on group fairness in social media, presenting a hierarchical schema for bias evaluation and a chain-of-thought mitigation method, GF-Think, to reduce bias. These contributions underline possible approaches to bias mitigation, aiming for ethical and equitable use of LLMs.
4.2. Worker Autonomy and Well Being
@@
Chhibber et al. (Chhibber et al., 2022) demonstrate how teachable conversational agents in crowd-based applications can empower workers, enhancing control, personalized learning, and ownership, thereby boosting job satisfaction. Konstantis et al. (Konstantis et al., 2023) emphasize transparency and fairness in crowdsourcing platforms, advocating for clear decision-making and respectful treatment to improve worker experience. Furthermore, Pal (Pal, 2023) highlights the design of LLMs focused on worker well-being and autonomy, avoiding stress-inducing features and promoting engagement and learning.
Despotovic and Bogodistov (Despotovic and Bogodistov, 2024) highlight a trend in job seekers preferring roles that incorporate advanced AI, like ChatGPT, aligning with their identities and technological interests. This shift towards AI-interactive jobs indicates a broader preference for technologically engaging roles, impacting job satisfaction and autonomy. Complementarily, Wang (Wang, 2023) explores the paradox of LLMs in the job market, noting their role in both creating new opportunities and obsoleting certain jobs. This dual effect is key in assessing worker autonomy and well-being, as LLM-driven automation fosters more creative roles and autonomy but also raises concerns about job security and satisfaction. Integrating insights from recent research, it’s evident that while generative AI has the potential to automate menial tasks and foster more creative roles, it also poses challenges to worker autonomy, necessitating a balanced approach to AI deployment in the workplace (Woodruff et al., 2023).
4.3. Effect on Wages and Jobs
@@
Hemmer et al. (Hemmer et al., 2023) suggest AI could enhance productivity by up to 40%, potentially leading to higher wages. In contrast, Chowdhury et al. (Chowdhury et al., 2022) point to AI’s role in both creating and displacing jobs, with an estimated 1.7 million new jobs and significant displacements in the U.S. by 2029. This highlights the necessity for proactive support for those affected by AI-driven changes. Chhibber et al. (Chhibber et al., 2022) discuss AI’s impact on the gig economy, noting potential job losses in areas like customer service, balanced by new opportunities in AI-related fields. Pal (Pal, 2023) foresees a growing LLM industry, suggesting increased job creation in AI R&D. Conversely, Wang (Wang, 2023) anticipates the displacement of up to 800 million jobs globally by 2030, advocating for protective legislation and equitable AI benefit distribution. Ashktorab et al. (Ashktorab et al., 2020) examine Human-AI collaboration in the gaming industry, where AI’s role in automating tasks like game testing could lead to job enrichment in creative roles.
@@
The integration of Large Language Models has diverse implications for the job market. Despotovic and Bogodistov’s (Despotovic and Bogodistov, 2024) research reveals that job seekers show a preference for roles that resonate with their personal identity and involve frequent interaction with technologies like ChatGPT. This sees a shift in job market dynamics and the necessity for businesses to adapt their recruitment strategies. Wang (Wang, 2023) discusses the paradoxical effect of LLMs, highlighting both job creation and displacement across various sectors. This includes an examination of LLMs’ influence on routine tasks, specialized professions, and broader economic and societal aspects. Walkowiak et al. (Walkowiak and MacDonald, 2023) focus on the Australian workforce, quantifying the exposure to GenAI risks and potential job transformations, addressing privacy, cybersecurity, and ethical issues.
4.4. Data Privacy and Security
@@
In the realm of human-AI collaboration, trust, particularly in data privacy and security, is foundational. Ezer et al. (Ezer et al., 2019) emphasize the critical role of trust engineering, advocating for algorithmic transparency and dynamic management to establish mutual trust in complex, information-imperfect environments. Yin et al. (Yin et al., 2019) further this by proposing a privacy-focused, human-centric approach, leveraging logistic regression and differential privacy to strike a balance between privacy and utility, achieving notable classification accuracy. In medical imaging, Kaissis et al. (Kaissis et al., 2020) highlight the imperative for secure, privacy-preserving AI through federated learning, addressing the privacy-utilization dichotomy to advance patient care. The synergy of AI and cybersecurity, as explored by Admin et al. (Samyuktha et al., 2022), showcases AI’s potential to fortify cyber defenses, underscoring the role of machine and deep learning in countering complex cyber threats and safeguarding data integrity. Lepri et al. (Lepri et al., 2021) advocate for a human-centric AI ethos, stressing the importance of privacy, accountability, and ethical decision-making in building secure, trustworthy AI systems.
@@
Regulatory and privacy challenges for Large Generative AI Models (LGAIMs), including ChatGPT and GPT-4, demand targeted legal and technical solutions for safe use. Hacker et al. (Hacker et al., 2023) recommend a regulatory framework for LGAIMs focusing on transparency, risk management, and equitable treatment, particularly for high-risk uses. Ullah et al. (Ullah et al., 2023) propose PrivChatGPT, a model incorporating differential privacy for LLMs, highlighting the balance between privacy and utility. Gupta et al. (Gupta et al., 2023) discuss cybersecurity risks with GenAI, emphasizing the need for ethical practices and secure defenses against potential misuse. De Angelis et al. (De Angelis et al., 2023) identify LLMs as vectors for misinformation, advocating for policies to combat the "AI-driven infodemic" in public health. Thapa and Adhikari (Thapa and Adhikari, 2023) balance LLMs’ benefits against misinformation risks in biomedical research, calling for strict validation. Sebastian (Sebastian, 2023) focuses on privacy in AI chatbots, suggesting federated learning and data minimization to protect user information, reflecting increasing privacy concerns among LLM users.
4.5. Trustworthy AI and Accountability
@@
Initially, conventional methods in trustworthy AI emphasized the importance of ethical principles, human oversight, and transparency. Hou et al. (Hou et al., 2023) discuss these principles, stressing the need for secure systems that identify risks while remaining under human control. Pflanzer et al. (Pflanzer et al., 2022) introduce the Agent-Deed-Consequence model as a traditional ethical algorithm framework, emphasizing human oversight in AI ethics. Caldwell et al. (Caldwell et al., 2022) highlight factors influencing trust in AI in cooperative environments, such as multiplayer gaming, underscoring the importance of environmental factors and peer influence in conventional trust-building.
@@
The advent of LLMs introduces new dimensions to these conventional methods. In healthcare, a case study using LLMs (Wang, 2023) showcases the balance between patient engagement and AI over-reliance risks, underscoring the advancements in data sensitivity and patient communication. Kim et al. (Kim et al., 2023) demonstrate how LLMs contribute to managing uncertainty in human-AI collaboration, proposing a trial-to-trial approach to effectively communicate and estimate uncertainty. Furthermore, LLMs influence the formation of mental models, as discussed by Bansal et al. (Bansal et al., 2019). They highlight LLMs’ role in enhancing predictability and understanding of AI’s error boundaries, crucial for developing accurate mental models for better human-AI collaboration.
4.6. Law and Public Policy
@@
Traditionally, AI legal and policy frameworks have focused on ensuring transparency, accountability, fairness, and bias mitigation. Pioneering work by Jobin et al. (Jobin et al., 2019), which analyzed AI guidelines globally, highlights the universal emphasis on ethical standards, transparency, and accountability. Similarly, research by Cath et al. (Cath et al., 2018) highlights the critical need for transparent and accountable AI, especially in decision-making processes. Stahl et al.(Stahl et al., 2021) propose six guidelines for AI regulation: Transparency, Accountability, Bias Mitigation, Human Oversight, Data Protection, and Public Education.
@@
The integration of Large Language Models (LLMs) into the AI landscape underscores the need for updated legal frameworks to address new challenges in data use and AI-generated content regulation. Bender et al. (Binz et al., 2023) examine the ethical implications of LLMs, focusing on data usage and biases. An accredited law partner (Walsh, 2023) specializing in technology and intellectual property discusses U.S. court cases like GitHub Copilot’s use of open-source code, highlighting the legal intricacies of data attribution and compliance. This situation prompts a reevaluation of the datasets used for AI training and the adaptation of laws, particularly concerning copyrights and patents. Ekenobi et al. (ThankGod Chinonso, 2023) explore the legal impacts of deploying LLMs, with a focus on data protection and privacy. They commend ChatGPT’s privacy measures as a step towards mitigating potential risks. Additionally, Marcos and Pullin’s article (Marcos and Pullin, 2023) talks about the EU’s GDPR challenges, emphasizing the balance between technological growth and privacy rights. This necessitates robust regulatory frameworks that ensure LLM compliance with strict data protection standards, highlighting the importance of new laws in safeguarding privacy while fostering innovation.
5. Applications